OpenStudy (alurahenderson):

WILL MEDAL The area of the trapezoid is 39 ft2. What is the height of the trapezoid?

OpenStudy (anonymous):

area=(a+b)/2*h 39=(15+11)/2 h 39*2=26h h=39*2/26

OpenStudy (alurahenderson):

3?